OpenStudy (girlgoyleh.):

What is the measure of each interior angle in a regular 15-gon? A. 120° B. 167° C. 108° D. 156°

OpenStudy (girlgoyleh.):

@mathstudent55

OpenStudy (jdoe0001):

\(\bf \textit{sum of interior angles in a regular polygon}={\color{brown}{ n}}\theta=180({\color{brown}{ n}}-2) \\ \quad \\ {\color{brown}{ n}}=\textit{sides in the polygon}\qquad \textit{solve for }\theta\impliedby \textit{interior angle}\)

OpenStudy (jdoe0001):

well.... hmmm \(\bf \textit{sum of interior angles in a regular polygon}\implies{\color{brown}{ n}}\theta=180({\color{brown}{ n}}-2) \\ \quad \\ {\color{brown}{ n}}=\textit{sides in the polygon}\qquad \textit{solve for }\theta\impliedby \textit{interior angle}\)

OpenStudy (girlgoyleh.):

so i mulitply 15 times 0?

OpenStudy (girlgoyleh.):

the equation would look like this 15x0=180(15-2)

OpenStudy (girlgoyleh.):

167?

OpenStudy (jdoe0001):

well.. . the 15-gon has 15 sides thus \(\bf \textit{sum of interior angles in a regular polygon} \\ \quad \\ {\color{brown}{ 15}}\theta=180({\color{brown}{ 15}}-2)\impliedby \textit{solve for }\theta\)

OpenStudy (girlgoyleh.):

B

OpenStudy (girlgoyleh.):

please hurry, my laptop is overheating and it shuts down by itself!!

OpenStudy (girlgoyleh.):

i think its B, am i correct?

OpenStudy (jdoe0001):

well.. what did you get for \(\theta?\)

OpenStudy (girlgoyleh.):

is that a zero? i multiplied 0 times 15 and then did the rest of the equation and got 167

OpenStudy (girlgoyleh.):

D